# Boundary Control and Calderón type Inverse Problems in Non-local heat equation

### v2

@@ Abstract.
We examine various density results related to the solutions of the non-local heat equation at a specific time slice, focusing on two distinct models: one with homogeneous Dirichlet boundary condition and the other with singular boundary data. In both the cases, we assume the non-local exponent $a\in(\frac{1}{2},1)$. We explore both the qualitative and quantitative aspects of the approximations. Additionally, we address Calderón-type inverse problems for these parabolic models, where we recover the potentials by analyzing the solutions either on the boundary or at a particular time slice. In both the density results and the Calderón type inverse problems, the Pohozaev identity plays a crucial role. Finally, in the last section, we apply the Pohozaev identity to a specific elliptic eigenvalue problem and demonstrate that the eigenfunctions, when divided by an appropriate power of the distance function, can not vanish on any non-empty open subset of the boundary. This particular eigenvalue problem does not need any restriction on the non-local exponent.
@@ 1. Introduction
In this article, we address various qualitative and quantitative approximation properties of solutions to certain non-local heat equation. Specifically, we consider two equations where one has homogeneous Dirichlet boundary condition and the other has singular boundary data. For the singular boundary data, we analyze the approximation properties by deriving an observability estimate through the Pohozaev identity [see [ROS14b]]. Using the same Pohozaev identity and density results, we also recover the potentials for both the systems by analyzing the solutions either at a specific time slice or on the boundary. We consider the following two equations:
(1) $\partial_{t}u_{f}+(-\Delta)^{a}u_{f}+qu_{f}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{f}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ u_{f}(0,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
@@
(2) $\partial_{t}u_{F}+(-\Delta)^{a}u_{F}+qu_{F}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{F}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{u_{F}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ u_{F}(0,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega.\end{aligned}\right.$
Here $\Omega\subset\mathbb{R}^{N}$ is a bounded $C^{1,1}$ domain with boundary $\partial\Omega$ and $a\in(\frac{1}{2},1)$. The functions $u_{f},u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ are the unknowns and the potential $q\in C_{c}^{\infty}\left((0,T)\times\Omega\right)$. Furthermore, for (1), we assume the initial condition $f\in C_{c}^{\infty}(\Omega)$ and for (2) we assume the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. The distance function ${d:\Omega\to\mathbb{R}_{\geq 0}}$ is the following:
@@
where $\|\cdot\|_{2}$ is the classical Euclidean norm in $\mathbb{R}^{N}$. To ensure uniqueness, we choose $q$ in such a way that zero is not an eigenvalue for either equation (1) or (2), corresponding to the cases where $f\equiv 0$ or $F\equiv 0$, respectively. Regarding approximation, in [RS17], the authors investigate a qualitative and quantitative approximation properties of solutions to a specific non-local heat equation, where they vary the $N+1$ Dirichlet data in the exterior domain. They showed that the solution space is dense in $\mathrm{L}^{2}$ in the interior by leveraging the unique continuation property: if a function $u$ satisfies $u=(-\Delta)^{a}u=0$ in some open set $W\subset\mathbb{R}^{N}$, then it must vanish identically, i.e., $u\equiv 0$ [GSU20]. The precise result is as follows:
@@
Let $a\in(0,1)$ and $B_{1}$ is the unit ball in $\mathbb{R}^{N}$ centered at origin. Let $W\subset\mathbb{R}^{N}$ be a non-empty open set such that $\overline{B}_{1}\cap\overline{W}$ is empty. Let for $f\in C_{c}^{\infty}((-1,1)\times W)$, the function $u_{f}:((-1,1)\times\mathbb{R}^{N})\to\mathbb{R}$ be the solution of the following equation:
(3) $\partial_{t}u_{f}+(-\Delta)^{a}u_{f}=&0\qquad\qquad\mbox{in}\ (-1,1)\times B_{1}\\ u_{f}=f&\qquad\qquad\ \mbox{in}\ (-1,1)\times\mathbb{R}^{n}\setminus B_{1}\\ u_{f}(-1,\cdot)=&f\qquad\qquad\mbox{in}\ B_{1}.\end{aligned}\right.$
@@
The quantitative analysis enables the authors to derive an estimate for the cost of approximation, as described in [RS17]. This requires the use of variational techniques, as presented in [RS17], [FCZ00], and [Lio92]. A qualitative version of similar approximation properties for the non-local elliptic model is presented in [GSU20], where the authors investigate the Calderón type problem for the fractional heat equation. A comparable result in the $C^{k}$ space is established in [DSV19], where the authors demonstrate that any $C^{k}$ function can be approximated by smooth solutions to the non-local heat equation, with the approximation property governed by the fractional component of the equation only. However, in this article, we vary either the $N$ dimensional initial data or boundary data, and obtain $\mathrm{L}^{2}$ approximation on an $N$-dimensional interior slice. In particular, we vary the initial data $f\in C_{c}^{\infty}(\Omega)$ for the equation with homogeneous Dirichlet boundary data (1), and the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ for the equation (2). We demonstrate that for a fix time slice $T>0$, the solution sets ${\{u_{f}:f\in C_{c}^{\infty}(\Omega)\}}$ and ${\{u_{F}:F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\}}$ corresponding to the equations (1) and (2) respectively, are dense in $\mathrm{L}^{2}(\Omega)$. The application of the integration by parts formula (in the $\mathrm{L}^{2}$ sense) imposes the following constraint $a\in(\frac{1}{2},1)$ on the non-local exponent. We think such result and techniques executed here could be extended to the classical heat equation i.e. $a=1$.
These results are partially motivated by the stability results for Calderón type inverse problems for the non-local heat equation, as studied in [RS17]. Our work can be seen as a continuation of the contributions made in [RS17] and [FCZ00], extending the analysis to boundary controllability in the presence of singular boundary data. Additionally, the study in [ROS14b] serves as a key motivation for exploring boundary controllability and Calderón-type inverse problems in the context of both singular boundary and initial data problem.
The density argument for the case of the singular boundary data (2) requires observability estimates, which we derive using the Pohozaev estimate, as demonstrated in [ROS14b]. Using density results and the same arguments presented in the proof of observability estimate, we solve the Calderón type inverse problems for both the non-local heat equations (1) and (2), where we analyze the solutions at a particular time slice $T>0$ for (1) and where we analyze the solutions on the boundary for (2). The study of the Calderón type inverse problem for the non-local elliptic equation can be found in [GLX17], [GRSU20], [Sal17], [Gho21]. The main results in this article are the following:
@@
Let $a\in(\frac{1}{2},1)$ and let $f\in C_{c}^{\infty}(\Omega)$ and let $u_{f}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (1). Then, for a fix time $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$.
@@
Let $a\in(\frac{1}{2},1)$ and let $F\in C_{c}^{\infty}((0,T)\times\Omega)$ and let $u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (2). Then, for a fix time $\tilde{T}>0$, the set
@@
(4) $\partial_{t}u_{i}+(-\Delta)^{a}u_{i}+q_{i}u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ u_{i}(0,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $q_{1},q_{2}\in C_{c}^{\infty}((0,T)\times\Omega)$ and $f\in C^{\infty}_{c}(\Omega)$. Furthermore, $q_{1},q_{2}$ be such that zero isn’t an eigenvalue for the same equation with zero initial data. If
@@
Let $a\in(\frac{1}{2},1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a-1(2a)}(\Omega))$ satisfy: for $i=1,2$
(5) $\partial_{t}u_{i}+(-\Delta)^{a}u_{i}+q_{i}u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{u_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ u_{i}(0,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $q_{1},q_{2}\in C_{c}^{\infty}((0,T)\times\Omega)$ and $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. Furthermore, $q_{1},q_{2}$ be such that zero isn’t an eigenvalue for (4) with zero boundary data. If
$u_{1}(T,\cdot)\equiv u_{2}(T,\cdot)\qquad\forall\,F\in C_{c}^{\infty}((0,T)\times\Omega)$
@@
In fact, we will show statements of theorem 4 and theorem 5 are equivalent. Next we describe solution spaces, well posedness of the solutions and various regularity results. Many of the cases, the notion of the solutions for non-local heat equation can be extended to non-local heat equation as the perturbation potential $q$ lies in $C_{c}^{\infty}((0,T)\times\Omega)$ [Paz12], [BJ12].
@@ 1.1. Preliminaries: Fractional Laplacian and fractional Sobolev space
Let $10<a<$ 1. The fractional Laplacian operator $(-\Delta)^{a}$ is defined over the space of Schwartz class functions $\mathcal{S}(\mathbb{R}^{N})$ as
@@
where $\hat{.}$ and $\mathcal{F}^{-1}$ denotes the Fourier and inverse Fourier transformation respectively. There are many equivalent definitions of fractional Laplacian [see [Kwa17]]. One of such equivalent definition we use throughout this article is given by the Cauchy principle value: for $0\leq a\leq 1$
@@
$H^{s}(\mathbb{R}^{N}):=\{u\in\mathbb{S}^{\prime}(\mathbb{R}^{N}):\langle\xi\rangle^{s}\hat{u}\in\mathrm{L}^{2}(\mathbb{R}^{N}).\},$
@@
$H^{s}(\Omega):=$$\{u=v|_{\Omega}:v\in H^{s}(\mathbb{R}^{N})\}.$
@@
$C^{s}_{c}(\mathbb{R}^{N}):=$$\{u\in C^{s}(\mathbb{R}^{N}):\,\mbox{supp}\,u\ \mbox{compactly contained in}\ \mathbb{R}^{N}\}.$
@@
$\|u\|_{H^{s}(\Omega)}=\|u\|_{H^{s}(\Omega)}=\|\langle\xi\rangle^{s}\hat{u}\|_{\mathrm{L}^{2}(\Omega)}.$
@@
$[u]_{C^{s}(\Omega)}:=\inf_{M\geq 0}\left\{M:|u(x)-u(y)|\leq M|x-y|^{s},\,\forall\,x,y\in\Omega\right\}.$
The spaces $C_{c}^{s}(\mathbb{R}^{N})$ and $C^{s}(\overline{\Omega})$ $(\Omega$ is bounded)))) can be endowed with the following norms respectively
@@
Note that $C^{s}(\overline{\Omega})$ is a Banach space with respect to the norm described abouve.
The definition of $C^{s}(\Omega)$ ((((or $C_{c}^{s}(\mathbb{R}^{N}))$ can be extended to any $\beta>0$. We can express $\beta=k+s$, where $k\in\mathbb{N}\cup\{0\}$ and $s\in(0,1)$. The space $C^{\beta}(\Omega)$ is the following

### v3

@@ Abstract.
We examine various density results related to the solutions of the non-local heat equation at a specific time slice, focusing on two distinct models: one with homogeneous Dirichlet boundary condition and the other with singular boundary data. In both the cases, we assume the non-local exponent $a\in(\frac{1}{2},1)$. We explore both the qualitative and quantitative aspects of the approximations. However, to obtain the approximation result for singular boundary data, we assume the potential to be small, non-negative, and exhibit only mild growth. The smallness of the potential is explicit in our study and depends on the domain and the dimension only. Additionally, we address Calderón-type inverse problems for these parabolic models, where we recover the potentials by analyzing the solutions either on the boundary or at a particular time slice. In both the density results and the Calderón type inverse problems, the Pohozaev identity plays a crucial role. Finally, in the last section, we apply the Pohozaev identity to a specific elliptic eigenvalue problem and demonstrate that the eigenfunctions, when divided by an appropriate power of the distance function, can not vanish on any non-empty open subset of the boundary. This particular eigenvalue problem does not need any restriction on the non-local exponent.
@@ 1. Introduction
In this article, we address various qualitative and quantitative approximation properties of solutions to certain non-local heat equation. Specifically, we consider two equations where one has homogeneous Dirichlet boundary condition and the other has singular boundary data. For the singular boundary data, we analyze the approximation properties through the Pohozaev identity [see [ROS14b]]. Using the same Pohozaev identity and density results, we also recover the potentials for both the systems by analyzing the solutions either at a specific time slice or on the boundary. We consider the following two equations:
(1) $\left\{\begin{aligned} \partial_{t}u_{f}+(-\Delta)^{a}u_{f}+qu_{f}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{f}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\Omega\\ u_{f}(0,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right$.
@@
(2) $\left\{\begin{aligned} \partial_{t}u_{F}+(-\Delta)^{a}u_{F}+qu_{F}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{F}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{u_{F}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ u_{F}(0,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega.\end{aligned}\right$.
Here $\Omega\subset\mathbb{R}^{N}$ is a bounded $C^{1,1}$ domain with boundary $\partial\Omega$ and $a\in(\frac{1}{2},1)$. The functions $u_{f},u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ are the unknowns and the potential $q\in C_{c}^{\infty}\left(\Omega\right)$. Furthermore, for (1), we assume the initial condition $f\in C_{c}^{\infty}(\Omega)$ and for (2) we assume the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. The distance function ${d:\Omega\to\mathbb{R}_{\geq 0}}$ is the following:
@@
where $\|\cdot\|_{2}$ is the classical Euclidean norm in $\mathbb{R}^{N}$. To ensure uniqueness, we choose $q$ such that zero is not an eigenvalue of either equation (1) or (2), corresponding to the cases where $f\equiv 0$ or $F\equiv 0$, respectively. Regarding approximation, in [RS17], the authors investigate a qualitative and quantitative approximation property of solutions to a specific non-local heat equation, where they vary the $N+1$ Dirichlet data in the exterior domain. They showed that the solution space is dense in $\mathrm{L}^{2}$ in the interior by leveraging the unique continuation property: if a function $u$ satisfies $u=(-\Delta)^{a}u=0$ in some open set $W\subset\mathbb{R}^{N}$, then it must vanish identically, i.e., $u\equiv 0$ [GSU20]. The precise result is as follows:
@@
Let $a\in(0,1)$ and $B_{1}$ is the unit ball in $\mathbb{R}^{N}$ centered at origin. Let $W\subset\mathbb{R}^{N}$ be a non-empty open set such that $\overline{B}_{1}\cap\overline{W}$ is empty. Let for $f\in C_{c}^{\infty}((-1,1)\times W)$, the function $u_{f}:(-1,1)\times\mathbb{R}^{N}\to\mathbb{R}$ be the solution of the following equation:
(3) $\left\{\begin{aligned} \partial_{t}u_{f}+(-\Delta)^{a}u_{f}=&0\qquad\qquad\mbox{in}\ (-1,1)\times B_{1}\\ u_{f}=&f\qquad\quad\ \ \ \mbox{in}\ (-1,1)\times\mathbb{R}^{N}\setminus B_{1}\\ u_{f}(-1,\cdot)=&f\qquad\qquad\mbox{in}\ B_{1}.\end{aligned}\right$.
@@
The quantitative analysis enables the authors to derive an estimate for the cost of approximation, as described in [RS17]. This requires the use of variational techniques, as presented in [RS17], [FCZ00], and [Lio92]. A qualitative version of similar approximation properties for the non-local elliptic model is presented in [GSU20], where the authors investigate the Calderón type problem for the fractional heat equation. A comparable result in the $C^{k}$ space is established in [DSV19], where the authors demonstrate that any $C^{k}$ function can be approximated by the smooth solutions to the non-local heat equation, with the approximation property governed by the fractional component of the equation only. However, in this article, we vary either the $N$ dimensional initial data or boundary data, and obtain $\mathrm{L}^{2}$ approximation on an $N$-dimensional interior slice. In particular, we vary the initial data $f\in C_{c}^{\infty}(\Omega)$ for the equation with homogeneous Dirichlet boundary data (1), and the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ for the equation (2). We demonstrate that for a fix time slice $T>0$, the solution sets ${\{u_{f}:f\in C_{c}^{\infty}(\Omega)\}}$ and ${\{u_{F}:F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\}}$ corresponding to the equations (1) and (2) respectively, are dense in $\mathrm{L}^{2}(\Omega)$. The application of the integration by parts formula (in the $\mathrm{L}^{2}$ sense) imposes the following constraint $a\in(\frac{1}{2},1)$ on the non-local exponent. We think such results and techniques executed here could be extended to the classical heat equation i.e., $a=1$.
These results are partially motivated by the stability results for Calderón type inverse problems for the non-local heat equation, as studied in [RS17]. Our work can be seen as a continuation of the contributions made in [RS17] and [FCZ00], extending the analysis to the boundary controllability in the presence of singular boundary data. Additionally, the study in [ROS14b] serves as a key motivation for exploring boundary controllability and Calderón type inverse problems in the context of both the singular boundary and the initial data problem.
The density argument for the case of the singular boundary data (2) requires the Pohozaev estimate, as demonstrated in [ROS14b]. For small potential, we can prove the density result using observability estimate as described in [BWZ25]. Using the density results and the same arguments presented in the proof of observability estimate, we solve the Calderón type inverse problems for both the non-local heat equations (1), and (2), where we analyze the solutions at a particular time slice $T>0$ for (1), and for (2) we analyze the solutions on the boundary. The study of the Calderón type inverse problems for the non-local elliptic equations can be found in [GLX17], [GRSU20], [Sal17], [Gho21]. The main results in this article are the following:
@@
Let $a\in(\frac{1}{2},1)$ and let $f\in C_{c}^{\infty}(\Omega)$. Let $u_{f}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (1). Then, for a fix time $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $\mathrm{L}^{2}(\Omega)$.
@@
Let $a\in(\frac{1}{2},1)$ and let $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. Let ${\theta<\frac{1}{2}\left(1+C_{HS}\left(\frac{N}{2}+R\right)\right)^{-1}}$, where $C_{HS}$ is the Hardy-Sobolev constant on the domain $\Omega$ and $R:=\max\{\|x\|,x\in\Omega\}$. Let the potential $q$ is non-negative and $|q(x)|,|\nabla q(x)|\leq\theta$ for all $x\in\Omega$. Let $u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (2). Then, for a fix time $\tilde{T}>0$, the set
@@
(4) $\left\{\begin{aligned} \partial_{t}u_{i}+(-\Delta)^{a}u_{i}+q_{i}u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\Omega\\ u_{i}(0,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right$.
where $q_{1},q_{2}\in C_{c}^{\infty}(\Omega)$ and $f\in C^{\infty}_{c}(\Omega)$. Furthermore, $q_{1},q_{2}$ be such that zero isn’t an eigenvalue to the same equation with zero initial data. If
@@
Let $a\in(\frac{1}{2},1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a-1(2a)}(\overline{\Omega}))$ satisfy: for $i=1,2$
(5) $\left\{\begin{aligned} \partial_{t}u_{i}+(-\Delta)^{a}u_{i}+q_{i}u_{i}=&0\qquad\qquad\ \mbox{in}\ (0,T)\times\Omega\\ u_{i}=&0\qquad\qquad\ \mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{u_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ u_{i}(0,\cdot)=&0\qquad\qquad\ \mbox{in}\ \Omega,\end{aligned}\right$.
where $q_{1},q_{2}\in C_{c}^{\infty}(\Omega)$ and $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. Furthermore, $q_{1},q_{2}$ be such that zero is not an eigenvalue to (4) with zero boundary data. If
$u_{1}(T,\cdot)\equiv u_{2}(T,\cdot)\qquad\forall\,F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$
@@
In fact, we will show that the statements of the theorem 4 and the theorem 5 are equivalent. Next we describe the solution spaces, well posedness of the solutions and various regularity results. Many of the cases, the notion of the solutions for the non-local heat equation can be extended to the non-local heat equation with the perturbation potential $q$, as the potential $q$ lies in $C_{c}^{\infty}(\Omega)$ [Paz12], [BJ12].
@@ 1.1. Preliminaries: Fractional Laplacian and fractional Sobolev space
Let $0<a<1$. The fractional Laplacian operator $(-\Delta)^{a}$ is defined over the space of Schwartz class functions $\mathcal{S}(\mathbb{R}^{N})$ as
@@
where $\hat{.}$ and $\mathcal{F}^{-1}$ denotes the Fourier and the inverse Fourier transformations respectively. There are many equivalent definitions of fractional Laplacian [see [Kwa17]]. One of such equivalent definition we use throughout this article is given by the Cauchy principle value: for $0\leq a\leq 1$
@@
$H^{s}(\mathbb{R}^{N}):=\{u\in\mathbb{S}^{\prime}(\mathbb{R}^{N}):\langle\xi\rangle^{s}\hat{u}\in\mathrm{L}^{2}(\mathbb{R}^{N})\},$
@@
$H^{s}(\Omega):=$ $\{u=v|_{\Omega}:v\in H^{s}(\mathbb{R}^{N})\}$
@@
$C^{s}_{c}(\mathbb{R}^{N}):=$ $\{u\in C^{s}(\mathbb{R}^{N}):\,\mbox{supp}\,u\ \mbox{compactly contained in}\ \mathbb{R}^{N}\},$
@@
$\|u\|_{H^{s}(\Omega)}=\|v\|_{H^{s}(\mathbb{R}^{N})}=\|\langle\xi\rangle^{s}\hat{v}\|_{\mathrm{L}^{2}(\mathbb{R}^{N})}.$
@@
$u]_{C^{s}(\Omega)}:=\inf_{M\geq 0}\left\{M:|u(x)-u(y)|\leq M|x-y|^{s},\,\forall\,x,y\in\Omega\right\}.$
The spaces $C_{c}^{s}(\mathbb{R}^{N})$ and $C^{s}(\overline{\Omega})$ $(\Omega$ is bounded)) can be endowed with the following norms respectively
@@
Note that $C^{s}(\overline{\Omega})$ is a Banach space with respect to the norm described above.
The definition of $C^{s}(\Omega)$ ((or $C_{c}^{s}(\mathbb{R}^{N}))$ can be extended to any $\beta>0$. We can express $\beta=k+s$, where $k\in\mathbb{N}\cup\{0\}$ and $s\in(0,1)$. The space $C^{\beta}(\Omega)$ is the following